Jess invested $30 more at 6% than at 5%. If the total annual interest is $100, how much is invested at each rate?
I've tried working this out but to know avail. I'm helping my son and I'm lost. Thanks

Answer by stanbon(75887)

Jess invested $30 more at 6% than at 5%. If the total annual interest is $100, how much is invested at each rate?
----------------------------
Equations:
Quantity Eq. : x = y + 30
Interest Eq. :0.06x + 0.05y = 100
------------------------------------------
Substitute:
0.06(y+30) + 0.05y = 100
Multiply thru by 100 to get:
6(y+30) + 5y = 10000
6y + 180 + 5y = 10000
11y = 9820
y = $892.73 (amount invested at 5%)
x = y + 30 = 892.73 + 30 = $922.73 (amount invested at 6%)
=============================================================
Cheers,
Stan H.